# Link Prediction with Untrained Message Passing Layers

### v1

@@ 4.2 Experimental Results
Results for attributed graphs are given in Table 1 where we find that the simplified model SGCN performs best on all attributed datasets, with performance comparable to more sophisticated state-of-the-art models Pan et al. (2021); Zhang and Chen (2018); Zhu et al. (2021) (See Table 9 in the Appendix) . Moreover, we find that in general simplified models perform better than or on par with their fully trained counterparts on almost all datasets, with the single exception of GIN on DBLP. We also find that the fully untrained (UT) architectures already provide a very good baseline and some cases even outperform fully trained versions. This demonstrates that the raw features produced by UTMP layers, which the simplified models are trained on, are already highly informative for link prediction in accordance with our theoretical results. Note that, the fully untrained (UT) models can be computed very efficiently via sparse matrix multiplication.
In the case of non-attributed graphs (Table 10) we observe that models based on UTMP layers achieve the highest score on 6 out of 8 datasets, with the exceptions being NS and Router datasets. Moreover, we find that the fully untrained UTGCN model outperforms all other models on the ’Celegans’, ’PB’, ’USAir’, ’E-coli’ which can be attributed to the reduced dimension of the learned features that come with the linear layers present in the simplified and fully trained models. In contrast to the trained variants UTGCN maintains the higher initial dimensionality of the features and hence can more efficiently discriminate between local neighborhoods by maintaining orthogonality. Furthermore, as we used one hot encodings as initial node features for the unattributed datasets orthonormality is satisfied exactly and therefore there is a one-to-one correspondence between node similarities produced by UTGCN and path based topological measures.
Table 2: Link Prediction accuracy for non-attributed networks as measured by ROC-AUC. Red values correspond to the overall best model for each dataset, and blue values indicate the best-performing model within the same category of message passing layers.
@@
Finally, we also examine the effect of increasing the number of UTMP layers using fully untrained (UT) models. Our results in Fig.1 indicate that, in general, UTGCN and UTSAGE maintain their performance as the number of layers is increased whereas the performance of UTGIN decreases sharply with more layers. This behavior can be attributed to the lack of degree based normalization in the formulation of GIN (see Sec.3.3) which leads UTGIN to be dominated by longer paths, and hence longer distance correlations, as the number of layers increases.
@@ 5 Conclusion
In this work, we explored the application of graph neural networks with untrained message passing layers for link prediction. Interestingly, our experimental evaluation of twelve data sets shows that simplifying GNNs architectures by eliminating trainable parameters and non-linearities not only enhances the link prediction performance of GNNs, but also improves their interpretability and training efficiency. As such untrained message passing layers offer a computationally efficient alternative to their fully trained counterparts that naturally scales to large graphs and in many cases produces state-of-the-art results. To complement our experimental results, we offered a theoretical perspective on untrained message passing, analytically establishing links between features generated by untrained message passing layers and path-based topological measures. We found that the link prediction offers a complementary perspective for analysing MPNNs and provides insights into the topological features captured by widely used initialization schemes such as random features and one-hot encodings.
@@
Acknowledgement
Lisi Qarkaxhija and Ingo Scholtes acknowledge funding from the German Federal Ministry of Education and Research (BMBF) via the Project "Software Campus 3.0", Grant No. (FKZ) 01IS24030, which is running from 01.04.2024 to 31.03.2026. Ingo Scholtes acknowledges funding through the Swiss National Science Foundation (SNF), Grant No. 176938.
@@ Appendix A Dataset details
Dataset ||||V|||| ||||E|||| Features Metric
Cora small Yang et al. [2016] 2,708 10,556 1,433 ROC-AUC
CiteSeer small Yang et al. [2016] 3,327 9,104 3,703 ROC-AUC
Cora Bojchevski and Günnemann [2017] 19,793 126,842 8,710 ROC-AUC
Cora ML Bojchevski and Günnemann [2017] 2,995 16,316 2,879 ROC-AUC
PubMed Bojchevski and Günnemann [2017] 19,717 88,648 500 ROC-AUC
CiteSeer Bojchevski and Günnemann [2017] 4,230 10,674 602 ROC-AUC
DBLP Bojchevski and Günnemann [2017] 17,716 105,734 1,639 ROC-AUC
NS Newman [2006] 1,461 2,742 - ROC-AUC
Celegans Watts and Strogatz [1998] 297 2,148 - ROC-AUC
PB Ackland et al. [2005] 1,222 16,714 - ROC-AUC
Power Watts and Strogatz [1998] 4,941 6,594 - ROC-AUC
Router Spring et al. [2004] 5,022 6,258 - ROC-AUC
USAir Batagelj and Mrvar [2006] 332 2,126 - ROC-AUC
Yeast Von Mering et al. [2002] 2,375 11,693 - ROC-AUC
E-coli Zhang et al. [2018] 1,805 15,660 - ROC-AUC
Appendix B Hyperparameter choices
@@
Appendix C Orthogonality of node features in empirical data sets
The distribution of inner products of initial node features for attributed datasets is given in Figure 2. We find that the inner products of feature vectors of randomly selected node pairs are in general close to zero. Note that, the feature vectors of all datasets are non-negative as they represent word occurrences. As expected, for connected nodes the inner products of feature vectors tend to be higher reflecting the increased feature similarity.
Appendix D Runtime Analysis and Training Efficency
@@
While we allocated a very generous limit of 10,000 epochs for training models in the main paper to ensure models can reach their best possible performance in order to compare the computational efficiency of the simplified models to their fully trained counterparts we also consider an experimental setting where we restrict the maximum number of training epochs to 100. We find that simplified models achieved convergence even for larger learning rates and considerably faster than their fully trained models. Even when constrained to 100 training epochs simplified models maintain scores that are almost identical to those presented in Table 8, while fully trained architectures suffer from the increased learning rates and require in general more epochs to converge. This leads to training efficiency gains similar to those reported by Wu et al. [2019] in the case of node classification.
@@
Cora small CiteSeer small Cora Cora ML PubMed CiteSeer DBLP
@@
Models Cora small CiteSeer small Cora Cora ML PubMed CiteSeer DBLP
@@
Appendix E Comparison to state-of-the-art
In Table 9 we compare UTMP architectures to three state-of-the-art link prediction methods, namely SEAL Zhang and Chen [2018], WalkPool Pan et al. [2021] and NBFNet Zhu et al. [2021] on attributed datasets. Compared to SEAL SGCN achieves higher scores on all 3 datasets. While NBFNet outperforms SGCN on Cora and PubMed by a small margin SGCN achieves a significantly higher score on CiteSeer. For WalkPoll the results on Cora and Citeseer are within one standard deviation of each other while WalkPool does better on PubMed.
For the unattributed datasets (see Table 10) we find that UTGCN has the highest overall score for E-coli while being within a standard deviation of the other methods on Celegans with WalkPool performing best on the remaining datasets. Note that WalkPool can be set up with different base methods as can be seen in Table 9 and hence using SGCN or UTGCN as a base method for WalkPool could potentially lead to performance improvements, which however is beyond the scope of this paper. It should also be noted that SGCN is considerably more efficient than all three methods, which rely either on subgraph extraction (SEAL and WalkPool) or learning explicit path representations (NBFNet).
Table 9: Link Prediction accuracy for attributed networks as measured by ROC-AUC compared to state-of-the-art models SEAL and NBFNet. ROC-AUC values for NBFNet and SEAL are from Zhu et al. [2021] (no standard deviations reported) and results for WalkPool from Pan et al. [2021].
Models Cora (small) CiteSeer (small) PubMed
SEAL 93.3 90.5 97.8
NBFNet 95.6 92.3 98.3
WalkPool (VGAE) 94.64 ± 0.55 94.32 ± 0.90 98.49 ± 0.13
WalkPool (ARGVA) 94.71 ± 0.85 94.53 ± 1.77 98.52 ± 0.14
WalkPool (GIC) 95.90 ± 0.50 95.94 ± 0.59 98.72 ± 0.10
GCN 92.82 ± 0.83 91.67 ± 1.14 97.54 ± 0.09
SGCN 95.3 ± 0.68 96.22 ± 0.4 97.94 ± 0.16
UTGCN 93.82 ± 0.68 96.0 ± 0.24 94.29 ± 0.24
Table 10: Link Prediction accuracy for non-attributed networks as measured by ROC-AUC compared to state-of-the-art models SEAL and WalkPool. ROC-AUC values SEAL are taken from Zhang and Chen [2018] and for WalkPool from Pan et al. [2021].
Models NS Celegans PB Power Router USAir Yeast E-coli
SEAL 97.71±0.93 89.54±2.04 95.01±0.34 84.18±1.82 95.68±1.22 97.09±0.70 97.20±0.64 97.22±0.28
WalkPool 98.95±0.41 92.79±1.09 95.60±0.37 92.56±0.60 97.27±0.28 98.68±0.48 98.37±0.25 98.58±0.19
[…]

### v2

@@ 4.2 Experimental Results
Results for attributed graphs are given in Table 1 where we find that in general replacing trained MP layers with their untrained counterparts increases LP performance on most datasets with the exception of PubMed and DBLP data sets where architectures based on UTMP layers perform worse in terms of Hits@100. In general, we find that GCN based architectures have the best overall performance. Finally, we observe that replacing trained GCN layers with their untrained counterpart also improves the LP performance of NCNC. Indeed, some of the results reported in [8] seem to be obtained using UTGCN layers. Although we focus is on MP layers rather end-to-end LP methods, we find that simple GAE type architectures based on UTMP layers can in many cases outperform more sophisticated state-of-the-art models such as SEAL, NBFnet and Neo-GNN (Sec. E).
We also find that the fully untrained (UT) architectures already provide a very good baseline and some cases even outperform fully trained versions in terms of ROC-AUC (Table 4). This demonstrates that the raw features produced by UTMP layers, which the simplified models are trained on, are already highly informative for link prediction in accordance with our theoretical results. Note that, the fully untrained (UT) models can be computed efficiently via sparse matrix multiplication.
In our analysis of the OGB datasets, we found that NCNC (GIN) outperforms other models in two out of three datasets. In the remaining dataset, SNCNC (SGIN) showed superior performance compared to the other models. Additionally, we observed that SNCNC models are highly competitive with fully trained models and, in some cases, are less memory-intensive. For instance, while the NCNC (GCN) model ran out of memory, the SNCNC (SGCN) model produced very good results without encountering this issue.
In the case of non-attributed graphs (Table 2) we observe that models based on UTMP layers achieve the highest score on 6 out of 8 datasets, with the exceptions being NS and Router datasets. Moreover, we find that the fully untrained UTGCN model performs best on the ’Celegans’, ’PB’, ’USAir’, ’E-coli’ which can be attributed to the reduced dimension of the learned features that come with the linear layers present in the simplified and fully trained models. Furthermore, as we used one hot encodings as initial node features for the unattributed datasets orthonormality is satisfied exactly and therefore there is a one-to-one correspondence between the UT models and path based topological measures.
Table 2: Link Prediction for non-attributed networks as measured by ROC-AUC.
@@
NCNC 92.66 ± 1.94 86.01 ± 3.13 95.27 ± 0.26 61.63 ± 2.18 73.06 ± 2.96 91.10 ± 2.14 93.41 ± 0.46 99.53 ± 0.09
SNCNC 91.28 ± 2.97 88.18 ± 2.65 95.77 ± 0.22 68.41 ± 1.46 87.29 ± 1.20 93.95 ± 1.36 95.42 ± 0.4 99.62 ± 0.05
Finally, we also examine the effect of increasing the number of UTMP layers using fully untrained (UT) models. Our results in Fig.2 indicate that, in general, UTGCN and UTSAGE maintain their performance as the number of layers is increased whereas the performance of UTGIN decreases sharply with more layers. This behavior can be attributed to the lack of degree based normalization in the formulation of GIN (see Sec.3.3) which leads UTGIN to be dominated by longer paths, and hence longer distance correlations, as the number of layers increases. In general however we find that UTMP layers do not suffer from over-squashing when equipped with proper degree based normalisation which can be attributed to the absence of nonlinearities and mixing between feature dimensions in UTMP layers.
@@ 5 Conclusion
In this work, we explored the application of graph neural networks with untrained message passing layers for link prediction. Interestingly, our experimental evaluation shows that simplifying GNN architectures by eliminating trainable parameters and nonlinearities can enhance link prediction performance and training efficiency. As such, untrained message passing layers offer a computationally efficient alternative to their fully trained counterparts that naturally scales to large graphs, while providing clearer theoretical links to classical path-based heuristics. To complement our experimental results, we offered a theoretical perspective on untrained message passing, analytically establishing links between features generated by untrained message passing layers and path-based topological measures. We found that the link prediction offers a complementary perspective for analysing MPNNs and provides insights into the topological features captured by widely used initialization schemes such as random features and one-hot encodings.
@@
Acknowledgements
Lisi Qarkaxhija and Ingo Scholtes acknowledge funding from the German Federal Ministry of Education and Research (BMBF) via the Project "Software Campus 3.0", Grant No. (FKZ) 16IS24030.
@@ Appendix A Dataset details
Summary statistics and sources of data sets are given in Table 3.
@@
Dataset ||V|| ||E|| Features
Cora small [45] 2,708 10,556 1,433
CiteSeer small [45] 3,327 9,104 3,703
Cora [46] 19,793 126,842 8,710
Cora ML [46] 2,995 16,316 2,879
PubMed [46] 19,717 88,648 500
CiteSeer [46] 4,230 10,674 602
DBLP [46] 17,716 105,734 1,639
OGBL-Collab [47] 235,868 1,285,465 128
OGBL-PPA [47] 576,289 30,326,273 58
OGBL-DDI [47] 4,267 1,334,889 -
NS [48] 1,461 2,742 -
Celegans [49] 297 2,148 -
PB [50] 1,222 16,714 -
Power [49] 4,941 6,594 -
Router [51] 5,022 6,258 -
USAir [52] 332 2,126 -
Yeast [53] 2,375 11,693 -
E-coli [54] 1,805 15,660 -
Appendix B Additional Results
In this section, we present additional results for the attributed datasets measured using ROC-AUC, and we also include the results for the OGB datasets.
ROC-AUC ROC-AUC ROC-AUC ROC-AUC ROC-AUC ROC-AUC ROC-AUC
SGIN 92.87 ± 0.37 93.6 ± 0.48 97.82 ± 0.11 95.26 ± 0.41 96.37 ± 0.23 94.13 ± 0.4 95.85 ± 0.11
Table 5: Link Prediction accuracy for OGB datasets. Red values correspond to the overall best model for each dataset, and blue values indicate the best-performing model within the same category of message passing layers.
Models ogbl-collab ogbl-ppa ogbl-ddi
Hits@50 Hits@100 Hits@20
GCN 55.58 ± 3.84 30.84 ± 1.78 48.58 ± 7.11
SGCN 54.13 ± 0.97 12.34 ± 1.95 22.14 ± 3.38
SAGE 45.90 ± 8.67 23.88 ± 1.63 22.00 ± 14.54
SSAGE 49.92 ± 1.52 8.86 ± 0.80 19.44 ± 10.84
NCNC (GCN) 62.29 ± 3.34 OOM 46.53 ± 29.24
SNCNC (SGCN) 65.40 ± 0.46 46.02 ± 1.19 13.55 ± 13.48
NCNC (SAGE) 64.91 ± 1.50 55.60 ± 3.06 40.38 ± 19.71
SNCNC (SSAGE) 62.31 ± 1.90 38.17 ± 2.37 13.19 ± 9.33
NCNC (GIN) 65.87 ± 0.74 59.47 ± 1.60 32.62 ± 30.71
SNCNC (SGIN) 16.69 ± 2.43 21.81 ± 0.75 54.53 ± 11.02
Appendix C Experimental setup and Hyperparameter choices
We use Adam [55] as an optimization function and employ binary cross entropy with logits as our loss function. All datasets are preprocessed by normalizing the node features and randomly splitting them. For non-attributed datasets, $10\%$ of the data is allocated to the test set, $5\%$ to the validation set [7], and the remaining data is used for the training set. In contrast, for attributed datasets, the split is $20\%$ for the test set, $10\%$ for the validation set, and the remainder for the training set [8].Each model configuration is run $10$ times, with the results averaged over these runs. Our training and testing procedures are based on the methodology outlined in [56], where we perform a new round of negative edge sampling for each training epoch. We limit the maximum number of epochs to 10,000 and also incorporate an early stopping mechanism in our training process by terminating training whenever there is no improvement in the validation set results over a span of $250$ epochs.
@@ Appendix C Experimental setup and Hyperparameter choices
Appendix D Runtime Analysis and Training Efficiency
@@
While we allocated a very generous limit of 10,000 epochs for training models in the main paper to ensure models can reach their best possible performance in order to compare the computational efficiency of the simplified models to their fully trained counterparts we also consider an experimental setting where we restrict the maximum number of training epochs to 100. We find that simplified models achieved convergence even for larger learning rates and considerably faster than their fully trained models. Even when constrained to 100 training epochs simplified models maintain scores that are almost identical to those presented in Table 10, while fully trained architectures suffer from the increased learning rates and require in general more epochs to converge. This leads to training efficiency gains similar to those reported by [9] in the case of node classification.
@@
Cora (small) CiteSeer (small) Cora Cora ML PubMed CiteSeer DBLP
@@
To complement our empirical runtime measurements, we provide an asymptotic analysis of the computational cost. The complexity per layer is as follows:
  • Trained GNN: $O(|\mathbf{E}|\cdot d\cdot h+|\mathbf{V}|\cdot h^{2})$, where the first term covers aggregation and the second term accounts for linear transformations.
  • UTMP: $O(|\mathbf{E}|\cdot d)$ for aggregation only.
Here, $d$ is the input feature dimension and $h$ is the hidden dimension. The key saving comes from UTMP eliminating the $O(|\mathbf{V}|\cdot h^{2})$ component related to learnable weight matrices, which is often the bottleneck in training GNNs on large graphs. This makes UTMP layers significantly more efficient, especially in resource-constrained settings. For simplified models (SMPNN), feature propagation can be precomputed once, reducing training to optimizing a single linear layer.
Appendix E HeaRT Split Results
In this section, we show the results of the datasets split by the HeaRT evaluation setting [44].
Table 11: Results on Cora, Citeseer, and Pubmed (%) under HeaRT.
Models Cora Citeseer Pubmed MRR Hits@10 MRR Hits@10 MRR Hits@10 Heuristic CN 9.78 20.11 8.42 18.68 2.28 4.78 AA 11.91 24.10 10.82 22.20 2.63 5.51 RA 11.81 24.48 10.84 22.86 2.47 4.9 Shortest Path 5.04 15.37 5.83 16.26 0.86 0.38 Katz 11.41 22.77 11.19 24.84 3.01 5.98 Embedding Node2Vec 14.47 ± 0.60 32.77 ± 1.29 21.17 ± 1.01 45.82 ± 2.01 3.94 ± 0.24 8.51 ± 0.77 MF 6.20 ± 1.42 15.26 ± 3.39 7.80 ± 0.79 16.72 ± 1.99 4.46 ± 0.32 9.42 ± 0.87 MLP 13.52 ± 0.65 31.01 ± 1.71 22.62 ± 0.55 48.02 ± 1.79 6.41 ± 0.25 15.04 ± 0.67 GNN GCN 16.61 ± 0.30 36.26 ± 1.14 21.09 ± 0.88 47.23 ± 1.88 7.13 ± 0.27 15.22 ± 0.57 GAT 13.84 ± 0.68 32.89 ± 1.27 19.58 ± 0.84 45.30 ± 1.3 4.95 ± 0.14 9.99 ± 0.64 SAGE 14.74 ± 0.69 34.65 ± 1.47 21.09 ± 1.15 48.75 ± 1.85 9.40 ± 0.70 20.54 ± 1.40 GAE 18.32 ± 0.41 37.95 ± 1.24 25.25 ± 0.82 49.65 ± 1.48 5.27 ± 0.25 10.50 ± 0.46 GNN+Pairwise Info SEAL 10.67 ± 3.46 24.27 ± 6.74 13.16 ± 1.66 27.37 ± 3.20 5.88 ± 0.53 12.47 ± 1.23 BUDDY 13.71 ± 0.59 30.40 ± 1.18 22.84 ± 0.36 48.35 ± 1.18 7.56 ± 0.18 16.78 ± 0.53 Neo-GNN 13.95 ± 0.39 31.27 ± 0.72 17.34 ± 0.84 41.74 ± 1.18 7.74 ± 0.30 17.88 ± 0.71 NCN 14.66 ± 0.95 35.14 ± 1.04 28.65 ± 1.21 53.41 ± 1.46 5.84 ± 0.22 13.22 ± 0.56 NCNC 14.98 ± 1.00 36.70 ± 1.57 24.10 ± 0.65 53.72 ± 0.97 8.58 ± 0.59 18.81 ± 1.16 NBFNet 13.56 ± 0.58 31.12 ± 0.75 14.29 ± 0.80 31.39 ± 1.34 >24h >24h PEG 15.73 ± 0.39 36.03 ± 0.75 21.01 ± 0.77 45.56 ± 1.38 4.4 ± 0.41 8.70 ± 1.26 GNN + GAE GCN 18.33 ± 0.32 37.29 ± 0.79 25.76 ± 0.68 50.44 ± 1.26 5.21 ± 0.27 10.55 ± 0.52 SAGE 14.34 ± 0.42 32.15 ± 1.32 21.35 ± 0.57 44.20 ± 0.98 4.12 ± 0.06 7.79 ± 0.24 GIN 12.81 ± 0.54 26.41 ± 1.21 16.29 ± 0.77 38.61 ± 1.01 3.64 ± 0.07 6.75 ± 0.11 GraphConv 12.48 ± 0.53 24.06 ± 1.30 13.36 ± 0.65 28.77 ± 1.14 3.38 ± 0.36 5.89 ± 0.88 S-GNN + GAE S-GCN 15.89 ± 0.26 34.27 ± 0.97 18.01 ± 2.59 43.78 ± 2.89 5.07 ± 0.24 9.95 ± 0.64 S-SAGE 13.59 ± 0.25 32.28 ± 0.84 14.45 ± 2.26 34.88 ± 3.38 3.04 ± 0.15 5.53 ± 0.52 S-GIN 12.87 ± 0.38 28.82 ± 1.30 12.89 ± 1.18 29.60 ± 2.78 3.78 ± 0.11 7.45 ± 0.15 NT-GNN + GAE NT-GCN 12.14 26.38 26.00 51.65 5.59 10.18 NT-SAGE 11.75 30.74 12.74 39.56 2.59 4.44 NT-GIN 0.77 0.19 1.16 1.10 1.89 1.94 NCNC GCN 15.77 ± 0.59 36.51 ± 0.98 21.43 ± 1.12 49.56 ± 1.74 9.54 ± 0.59 21.84 ± 1.00 SAGE 12.18 ± 0.57 30.34 ± 1.25 16.22 ± 1.72 31.34 ± 4.09 6.48 ± 0.37 14.84 ± 1.10 GIN 12.23 ± 0.48 30.76 ± 1.17 16.77 ± 1.05 33.43 ± 2.64 6.54 ± 0.27 14.62 ± 0.46 NCNC SGCN 15.92 ± 0.80 38.04 ± 0.63 23.55 ± 0.82 53.72 ± 1.34 8.29 ± 0.65 18.28 ± 1.49 SSAGE 11.92 ± 1.35 29.96 ± 0.93 20.04 ± 3.61 42.90 ± 6.60 4.87 ± 0.66 10.41 ± 1.57 SGIN 12.59 ± 0.75 30.06 ± 2.22 22.35 ± 0.83 45.87 ± 1.72 4.95 ± 0.30 10.38 ± 0.70
Table 12: Results on OGB datasets (%) under HeaRT.
Models ogbl-collab ogbl-ddi ogbl-ppa ogbl-citation2 MRR Hits@20 MRR Hits@20 MRR Hits@20 MRR Hits@20 CN 4.20 16.46 6.71 38.69 25.70 68.25 17.11 41.73 AA 5.07 19.59 6.97 39.75 26.85 70.22 17.83 43.12 RA 6.29 24.29 8.70 44.01 28.34 71.50 17.79 43.34 Shortest Path 2.66 15.98 0 0 0.54 1.31 >24h >24h Katz 6.31 24.34 6.71 38.69 25.70 68.25 14.10 35.55 Node2Vec 4.68 ± 0.08 16.84 ± 0.17 11.14 ± 0.95 63.63 ± 2.05 18.33 ± 0.10 53.42 ± 0.11 14.67 ± 0.18 42.68 ± 0.20 MF 4.89 ± 0.25 18.86 ± 0.40 13.99 ± 0.47 59.50 ± 1.68 22.47 ± 1.53 70.71 ± 4.82 8.72 ± 2.60 29.64 ± 7.30 MLP 5.37 ± 0.14 16.15 ± 0.27 N/A N/A 0.98 ± 0.00 1.47 ± 0.00 16.32 ± 0.07 43.15 ± 0.10 GCN 6.09 ± 0.38 22.48 ± 0.81 13.46 ± 0.34 64.76 ± 1.45 26.94 ± 0.48 68.38 ± 0.73 19.98 ± 0.35 51.72 ± 0.46 GAT 4.18 ± 0.33 18.30 ± 1.42 12.92 ± 0.39 66.83 ± 2.23 OOM OOM OOM OOM SAGE 5.53 ± 0.5 21.26 ± 1.32 12.60 ± 0.72 67.19 ± 1.18 27.27 ± 0.30 69.49 ± 0.43 22.05 ± 0.12 53.13 ± 0.15 GAE OOM OOM 3.49 ± 1.73 17.81 ± 9.80 OOM OOM OOM OOM SEAL 6.43 ± 0.32 21.57 ± 0.38 9.99 ± 0.90 49.74 ± 2.39 29.71 ± 0.71 76.77 ± 0.94 20.60 ± 1.28 48.62 ± 1.93 BUDDY 5.67 ± 0.36 23.35 ± 0.73 12.43 ± 0.50 58.71 ± 1.63 27.70 ± 0.33 71.50 ± 0.68 19.17 ± 0.20 47.81 ± 0.37 Neo-GNN 5.23 ± 0.9 21.03 ± 3.39 10.86 ± 2.16 51.94 ± 10.33 21.68 ± 1.14 64.81 ± 2.26 16.12 ± 0.25 43.17 ± 0.53 NCN 5.09 ± 0.38 20.84 ± 1.31 12.86 ± 0.78 65.82 ± 2.66 35.06 ± 0.26 81.89 ± 0.31 23.35 ± 0.28 53.76 ± 0.20 NCNC 4.73 ± 0.86 20.49 ± 3.97 >24h >24h 33.52 ± 0.26 82.24 ± 0.40 19.61 ± 0.54 51.69 ± 1.48 NBFNet OOM OOM >24h >24h OOM OOM OOM OOM PEG 4.83 ± 0.21 18.29 ± 1.06 12.05 ± 1.14 50.12 ± 6.55 OOM OOM OOM OOM GCN + GAE OOM OOM 6.52 ± 0.51 34.48 ± 1.52 OOM OOM OOM OOM SAGE + GAE OOM OOM 5.33 ± 0.21 29.30 ± 1.44 OOM OOM OOM OOM GIN + GAE OOM OOM 12.15 ± 0.26 54.21 ± 0.95 OOM OOM OOM OOM GraphConv + GAE OOM OOM 13.60 ± 0.26 56.36 ± 0.45 OOM OOM OOM OOM SGCN + GAE OOM OOM 3.82 ± 1.97 18.21 ± 10.38 OOM OOM OOM OOM SSAGE + GAE OOM OOM 2.71 ± 0.22 10.98 ± 0.76 OOM OOM OOM OOM SGIN + GAE OOM OOM 0.40 0.00 OOM OOM OOM OOM NTGCN + GAE 1.07 0.66 5.65 33.23 19.58 30.60 0.40 0.00 NTSAGE + GAE 0.40 0.00 0.60 0.00 1.96 3.26 0.40 0.00 NTGIN + GAE 0.40 0.00 0.40 0.00 0.40 0.00 0.40 0.00 NCNC + GCN 3.53 ± 0.29 13.99 ± 1.44 OOM OOM 35.75 ± 1.23 75.94 ± 1.51 21.52 ± 0.71 53.08 ± 1.61 NCNC + SAGE 2.97 ± 0.63 11.28 ± 2.77 OOM OOM 35.37 ± 1.09 71.62 ± 1.70 20.73 ± 0.63 51.87 ± 0.86 NCNC + GIN 2.93 ± 0.64 11.28 ± 2.56 OOM OOM 18.81 ± 6.81 38.51 ± 9.97 21.36 ± 0.89 52.48 ± 1.09 NCNC + SGCN 3.60 ± 0.87 14.06 ± 3.79 OOM OOM 33.98 ± 1.90 70.87 ± 3.55 20.25 ± 0.38 50.95 ± 0.33 NCNC + SSAGE 3.79 ± 0.54 15.17 ± 2.15 OOM OOM 34.09 ± 0.55 69.77 ± 1.57 20.55 ± 0.29 51.08 ± 0.39 NCNC + SGIN 2.38 ± 0.25 6.72 ± 0.85 OOM OOM 20.46 ± 1.41 40.16 ± 1.90 20.97 ± 0.86 51.13 ± 1.70